# FireFly-S: Exploiting Dual-Side Sparsity for Spiking Neural Networks Acceleration with Reconfigurable Spatial Architecture

### v1

@@ Abstract
Spiking Neural Networks (SNNs), with their brain-inspired structure using discrete spikes instead of continuous activations, are gaining attention for their potential of efficient processing on neuromorphic chips. While current SNN hardware accelerators often prioritize temporal spike sparsity, exploiting sparse synaptic weights offers significant untapped potential for even greater efficiency. To address this, we propose FireFly-S, a Sparse extension of the FireFly series. This co-optimized software-hardware design focusing on leveraging dual-side sparsity for acceleration. On the software side, we propose a novel algorithmic optimization framework that combines gradient rewiring for pruning and modified Learned Step Size Quantization (LSQ) tailored for SNNs, which achieves remarkable weight sparsity exceeding 85% and enables efficient 4-bit quantization with negligible accuracy loss. On the hardware side, we present an efficient dual-side sparsity detector employing a Bitmap-based sparse decoding logic to pinpoint the positions of non-zero weights and input spikes. The logic allows for the direct bypassing of redundant computations, thereby enhancing computational efficiency. Different from the overlay architecture adopted by previous FireFly series, we adopt a spatial architecture with inter-layer pipelining that can fully exploit the nature of Field-Programmable Gate Arrays (FPGAs). A spatial-temporal dataflow is also proposed to support such inter-layer pipelining and avoid long-term temporal dependencies. In experiments conducted on the MNIST, DVS-Gesture and CIFAR-10 datasets, the FireFly-S model achieves 85-95% sparsity with 4-bit quantization and the hardware accelerator effectively leverages the dual-side sparsity, delivering outstanding performance metrics of 10,047 FPS/W on MNIST, 3,683 FPS/W on DVS-Gesture, and 2,327 FPS/W on CIFAR-10.
@@ II-B Spiking Neural Networks Compression
As previously mentioned, SNNs are inherently sparse and event-driven, making them well-suited for efficient processing on neuromorphic chips. However, deploying SNNs on hardware platforms presents challenges due to their temporal dynamics and the memory requirements for storing membrane potentials. To overcome these challenges, researchers have explored various compression techniques, such as weight pruning and quantization, to further reduce parameter size and computational complexity in SNNs.
Accordingly, Compressing SNNs through sparsity regularization and ADMM optimization was proposed by Deng et al. [18]. Chen et al. [19] introduced a gradient-based rewiring method, jointly optimizing weight values and connections for pruning. Putra et al. [20] developed the Q-SpiNN framework, which quantizes SNNs by exploring various quantization schemes and precision levels, guided by a reward function. Wei et al. [21] propose a Weight-Spike Dual Regulation method to mitigate the performance degradation caused by quantization.
@@ III-A Compression Sequence
The sequence of pruning and quantization significantly influence the final performance of SNNs. Quantization is typically performed before pruning to preserve the network’s accuracy, as quantization ensures more paths for optimization before the connectivity is reduced by pruning. However, pruning can also affect the distribution of weights, potentially making the quantization scale and zero-point inappropriate for the pruned network. This can lead to suboptimal performance or require recalibration of the quantization parameters to maintain accuracy.
Another strategy is to perform pruning and quantization simultaneously, jointly optimizing the weight values and connections. This approach can explore a larger search space and potentially find better solutions, especially when pruning and quantization are integrated into the training process, which also avoids the need for additional post-training compression steps, retraining, or fine-tuning. This is precisely the strategy employed by FireFly-S, as shown in the Fig. 2 enabling it to achieve a favorable balance between compression and accuracy.
@@ III-B Compression Methodologies
The pruning algorithm we adopt is based on gradient rewiring [19]. In this method, each synaptic weight $w$ is represented as $w=s\odot\text{ReLU}(\theta)$, where $\theta$ represents the synaptic parameter, and $\odot$ denotes the element-wise product. The sign factor $s$, determined by the sign of the initial weight, indicates whether the connection is excitatory or inhibitory. This representation is illustrated in the left part of Fig. 2. During pruning, the optimization focuses on driving some $\theta$ values towards non-positive values, effectively eliminating the corresponding connections. Our approach extends this method by incorporating weight decay $\lambda$, which penalizes the magnitude of $\theta$ during optimization, driving more weights towards zero. Besides, to enable more comprehensive network compression, we apply this pruning strategy to both convolutional and fully connected layers, including bias parameters that optimized by AdamW [22].
@@ 9                   $A\leftarrow A\cup\{c_{o}\}$;
Motivated by the success of STE in approximating gradients for quantization, we present the LSQ variant specifically designed for SNNs. Unlike traditional LSQ, our approach jointly optimizes the quantization scale across weights, biases, and neuronal thresholds during training. We employ a shared scale for weights within the same channel, as well as for the corresponding bias and neuronal threshold. Additionally, we allow a wider quantization range for biases and thresholds and enable gradient descent on the neuronal thresholds. This facilitates end-to-end learning of quantized SNNs, as illustrated by the inference and training processes of a single layer depicted in the right part of Fig. 2. During training, weights, biases, and threshold values undergo quantization and dequantization, with the dequantized parameters used for forward and backward propagation. Conversely, inference relies solely on quantized parameters for forward propagation, eliminating the need for dequantization, scale storage, and associated computations.
@@ V Hardware Micro Architecture
The hardware microarchitecture of the FireFly-S is intricately designed to harness the dual-side sparsity characteristic of SNNs. It features a spatial layout that incorporates inter-layer pipelining, as illustrated in Fig. 5. Each pipeline stage is tailored to process a single layer of inference, equipped with specific modules including Padding, Dataflow Orchestrator, Dual-side Sparsity Detector, and Maxpooling.
Figure 5: Inter-layer Pipelined Architecture of FireFly-S. Each pipeline stage performs one layer inference and includes modules for Padding, Dataflow Orchestration, Dual-side Sparsity Detection, and Maxpooling.
The Padding module dynamically adjusts input dimensions for convolutional layers, adding padding in ”same” mode to match output and input sizes, while bypassing padding for ”valid” mode or fully connected layers. The Dataflow Orchestrator is pivotal in ensuring consistent dataflow between layers to maintain pipeline efficiency. It also efficiently manages the im2col operation for SNNs, essential for reorganizing input data into a format suitable for convolution operations, all while maintaining a membrane potential storage-free approach. The Dual-side Sparsity Detector module optimizes computations by identifying matching nonzero elements in both activation spikes and synaptic weights, enabling computational savings by bypassing unnecessary operations. Then, it generates specific read commands for relevant data (weights, biases, masks, thresholds) and efficiently updates neuron dynamics in the IF/LIF neuron model. Finally, the optional Maxpooling module provides spatial downsampling of the output, the output stream is then pipelined and fed into the next core or layer for further processing.
@@ V-A Dataflow Orchestrator
To reconcile these conflicting dataflow requirements and manage the im2col operation, FireFly-S employs a dedicated Dataflow Orchestrator, as illustrated in Fig. 6. This FIFO-based logic utilizes on-chip RAM divided into two sectors. The first sector functions as a minimal line buffer specifically for the im2col operation, storing only the data necessary to maintain the convolution window as the input feature map streams in. The size of this sector is determined by the input dimension and kernel size. Crucially, data within this sector is reused multiple times, as dictated by the parallelism level and data dimensions, to ensure continuous operation when parallelism cannot fully unroll the convolution. The remaining on-chip RAM serves as a configurable holding area, acting as a buffer for incoming spike feature maps from upstream processing. This holding area prevents pipeline stalls that would otherwise occur when the reused sector is full. Additionally, it provides a transition buffer for switching convolution windows. For clarity, the example in Fig. 6 depicts a parallelism level of 1, with an input dataflow of $(F_{h_{i}},F_{w_{i}},C_{i},T)$ being transformed to the output dataflow of $(F_{h_{i}},F_{w_{i}},C_{o},T,K_{h},K_{w},C_{i})$.
@@
The orchestrator leverages a combination of status registers, implemented as simple counters, and some stride counters. The status registers, chained to form loop iterators, pinpoint the current position within the dataflow. These registers don’t directly determine memory addresses but provide signals that guide the stride counters in generating the correct addresses for fetching data from the RAM. The use of stride counters for base address and offset address, each with its own specific increment value, allows the orchestrator to navigate the non-sequential memory access patterns required by the implicit im2col and data transpositions, eliminating the need for computationally expensive multiplication operations.
@@
The orchestrator manages data access through a push pointer and a pop pointer, mirroring the functionality of a traditional FIFO buffer. While the push pointer logic mirrors that of a standard FIFO, the pop pointer is determined by the combined values of the base and offset counters. The base counter marks the starting position of the current convolution window within the RAM, while the offset counter traverses elements within that window.
@@
For ”valid” convolution with stride 1, the base counter increment for the dataflow coordinator adapts based on the window’s position within the feature map. Under normal operation, the increment is simply $C_{i}\times T$, reflecting the number of data elements in the previous window’s channel and temporal dimensions. In addition to this regular increment, specific adjustments are needed for efficient transitions between rows and feature maps. After processing the last window of a row, the base counter must jump to the beginning of the next row, requiring an increment of $(K_{w}\times C_{i}\times T)$ to bypass the remaining $(K_{w}-1)$ columns of the previous row. Similarly, transitioning to a new feature map necessitates an increment of $((K_{h}-1)\times F_{w_{i}}+K_{w})\times C_{i}\times T$ to address the start of the new feature map’s data. This ability to directly address and access data within the RAM, regardless of its position, contributes to the efficiency of Dataflow Orchestrator. Traditional line buffers, reliant on sequential shifting, would inevitably incur overhead during both row and feature map transitions, as they must sequentially shift data to align the buffer with the required starting position for processing.
@@
Besides its role in managing the dataflow, the design logic of Dataflow Orchestrator also leverages the regularity of data access patterns, which are dictated by consistent stride patterns, to pre-calculate base and offset counter values. This pre-calculation enables pipelined address generation for the FIFO’s pop pointer, efficiently determining the next data element for processing and significantly enhancing throughput. Furthermore, the predictability arising from these consistent strides allows the FIFO’s full and empty signals to be implemented as registers. This design choice minimizes combinational logic delays, improving timing closure and ultimately contributing to support the  [^3] MHz working frequency across our entire architecture in our experiment, as discussed in Section VI.
@@ V-B Dual-side Sparsity Detector
[…]
The lower left corner of the Fig. 7 illustrates the initial transformation of the original sparse weight matrix into a Bitmap format. This process involves evaluating each matrix element to ascertain its non-zero status, resulting in the creation of a bitmap mask alongside a separate collection of non-zero values. These bitmaps are then segmented into vectors, each sized according to the input channel parallelism, $P_{C_{i}}$. A higher $P_{C_{i}}$ allows for larger bitmap segments, amplifying the throughput-to-storage advantage inherent in Bitmap representation.
@@ VI-A Experiments Setup
To evaluate the performance of FireFly-S, we conduct a series of experiments using various SNN models, specifically SCNN5, SCNN7, and SCNN9 models with 5, 7, and 9 convolutional layers. Detailed configurations of these models are presented in Table II, where ’c3p1’ indicates a 3x3 convolution kernel with padding of 1, ’mp2’ represents max pooling with a 2x2 filter and stride of 2, and ’10fc’ represents a fully connected layer with 10 output neurons. These models are trained on the MNIST, DVS-Gesture, and CIFAR-10 datasets, respectively. For testing, all models are deployed on the Xilinx Zynq Ultrascale+ FPGA, utilizing the KV260 Vision AI Starter Kit, which is specifically designed for the deployment of edge-based applications.

### v2

@@ Abstract
Spiking Neural Networks (SNNs), with their brain-inspired structure using discrete spikes instead of continuous activations, are gaining attention for their potential of efficient processing on neuromorphic chips. While current SNN hardware accelerators often prioritize temporal spike sparsity, exploiting sparse synaptic weights offers significant untapped potential for even greater efficiency. To address this, we propose FireFly-S, a Sparse extension of the FireFly series. This co-optimized software-hardware design focusing on leveraging dual-side sparsity for acceleration. On the software side, we propose a novel algorithmic optimization framework that combines gradient rewiring for pruning and modified Learned Step Size Quantization (LSQ) tailored for SNNs, which achieves remarkable weight sparsity exceeding 85% and enables efficient 4-bit quantization with negligible accuracy loss. On the hardware side, we present an efficient dual-side sparsity detector employing a Bitmap-based sparse decoding logic to pinpoint the positions of non-zero weights and input spikes. The logic allows for the direct bypassing of redundant computations, thereby enhancing computational efficiency. Different from the overlay architecture adopted by previous FireFly series, we adopt a parametric spatial architecture with inter-layer pipelining that can fully exploit the fine-grained programmability and reconfigurability of Field-Programmable Gate Arrays (FPGAs), enabling fast deployment for various models. A spatial-temporal dataflow is also proposed to support such inter-layer pipelining and avoid long-term temporal dependencies. In experiments conducted on the MNIST, DVS-Gesture and CIFAR-10 datasets, the FireFly-S model achieves 85-95% sparsity with 4-bit quantization and the hardware accelerator effectively leverages the dual-side sparsity, delivering outstanding performance metrics of 10,047 FPS/W on MNIST, 3,683 FPS/W on DVS-Gesture, and 2,327 FPS/W on CIFAR-10.
@@ II-B Spiking Neural Networks Compression
As previously mentioned, SNNs are inherently sparse and event-driven, making them well-suited for efficient processing on neuromorphic chips. However, hardware deployment is challenging due to increased computational complexity from temporal dynamics and a larger memory footprint from storing membrane potentials. To address these issues, compression techniques such as pruning and quantization have been explored to reduce the memory usage, and computational overhead, further enhancing the advantages of SNNs.
Accordingly, Compressing SNNs through sparsity regularization and ADMM optimization was proposed by Deng et al. [18]. Chen et al. [19] introduced a gradient-based rewiring method, jointly optimizing weight values and connections for pruning. These pruning techniques provide two significant benefits for hardware implementation. 1) Lower memory footprint. The reduction of synaptic connections can lead to fewer spikes and inactive neurons, which decreases memory access and weight storage. Additionally, the reduced input current allows for a smaller bitwidth for membrane potentials. 2) Enhanced computational efficiency. Fewer operations are needed for pruned connections, reducing complexity in proportion to the sparsity. For a network with 75% sparsity, the dense method processes 75% of computations that are ineffectual on average. However, by incorporating sparsity detection logic, the same performance can be achieved with just a quarter of the processing elements, eliminating wasted computations.
For quantization, Putra et al. [20] developed the Q-SpiNN framework, which quantizes SNNs by exploring various quantization schemes and precision levels, guided by a reward function. Wei et al. [21] propose a Weight-Spike Dual Regulation method to mitigate the performance degradation caused by quantization. By reducing the bitwidth, quantization not only lowers the memory footprint but also directly addresses the computational overhead challenges in neuromorphic hardware by simplifying the arithmetic operations.
@@ III-A Compression Sequence
The sequence of pruning and quantization shows impact on the performance of neural networks. Some studies advocate for a quantize-then-prune approach [22] [23], while others, such as Zhang et al. [24], suggest that the optimal sequence may depend on the network architecture. This indicates that different networks exhibit varying tendencies regarding the order of compression techniques. Therefore, treating pruning and quantization as separate steps may lead to suboptimal performance across various network, and any necessary post-processing or recalibration [25] [26] to maintain accuracy can become complex and challenging.
An alternative strategy is to jointly optimize pruning and quantization, where weight values and connections are simultaneously refined. This integrated approach explores a broader search space and has the potential to yield more optimal solutions [27] [28]. By embedding pruning and quantization within the training process, this strategy not only enhances the balance between compression and accuracy but also eliminates the need for additional post-training steps, such as retraining or fine-tuning. This joint optimization is precisely the strategy employed by FireFly-S, as depicted in Fig. 2.
@@ III-B Compression Methodologies
The pruning algorithm we adopt is based on gradient rewiring [19]. In this method, each synaptic weight $w$ is represented as $w=s\odot\text{ReLU}(\theta)$, where $\theta$ represents the synaptic parameter, and $\odot$ denotes the element-wise product. The sign factor $s$, determined by the sign of the initial weight, indicates whether the connection is excitatory or inhibitory. This representation is illustrated in the left part of Fig. 2. During training, $\theta$ is influenced by a Laplace prior distribution to impose a regulartization effect. Specifically, a penalty term $-\alpha\,\text{sign}(\theta-\mu)$ is applied, where $\alpha$ and $\mu$ represent the scale and location parameters, respectively[19].This penalty pushes some $\theta$ values towards non-positive numbers, thereby eliminating the corresponding connections.
Our approach builds on this by applying weight decay to $\theta$, scaling $\theta$ by $(1-\eta\lambda)$ during each update, where $\eta$ is the learning rate and $\lambda$ as the weight decay factor. This introduces an $\ell_{2}$ regularization effect, which not only encourages more $\theta$ values to approach zero, promoting sparsity, but also helps avoid sharper minima and improves generalization during optimization [29] [30]. Besides, to enable more comprehensive network compression, we apply this pruning strategy to both convolutional and fully connected layers, including bias parameters that optimized by AdamW [31].
@@ 9                   $A\leftarrow A\cup\{c_{o}\}$;
Building on the success of STE in approximating gradients for quantization, we introduce a novel LSQ variant tailored specifically for SNNs. The right part of Fig. 2 illustrates the quantization process in FireFly-S. During training, weights, biases, and thresholds are quantized and dequantized, with the dequantized values used for forward and backward propagation. In contrast, inference uses only quantized parameters for forward propagation, removing the need for dequantization, scale storage, and related computations.
Our approach differs from existing methods in two key ways: 1) We employ learnable per-channel scales, rather than relying on statistical methods like min-max or percentile-based approaches, which are commonly used in frameworks like SyncNN and earlier versions of the FireFly series. 2) Unlike standard LSQ and quantization frameworks like TensorRT, which often neglect bias quantization, our method jointly optimizes shared quantization scales for weights, biases, and neuronal thresholds within each channel during training.
These two key distinctions also offer two major benefits: 1) Learnable per-channel scales enable the model to capture and address quantization errors arising from various computations during training, including neuron dynamics like the decay process. This results in improved accuracy, as demonstrated in Section VI. This end-to-end training process ensures that, once trained, the model can be deployed directly without any additional precision loss. In contrast, statistical methods require a post-training process, and since the scales are based on weight statistics, applying them directly to biases and neuronal thresholds can cause mismatches and lead to accuracy degradation. 2) By sharing the quantization scale across weights within the same channel, along with the corresponding bias and neuronal threshold, our approach ensures consistent scaling across parameters. This eliminates the need for rescaling during weight accumulation, bias addition, and threshold comparison, enabling efficient integer-only inference. Additionally, due to the unique properties of spiking neurons, there is no need to rescale per-channel activations during inference compared to ANNs quantization, as the per-channel quantization scales are absorbed into the membrane thresholds. Based on these characteristics, our hardware accelerator can perform all computations using integer arithmetic, simplifying the computational logic and memory footprint.
@@ V Hardware Micro Architecture
The hardware microarchitecture of the FireFly-S is intricately designed to harness the dual-side sparsity characteristic of SNNs. It features a spatial layout that incorporates inter-layer pipelining, as illustrated in Fig. 5. Each pipeline stage is specifically designed to handle a single layer of inference within the FPGA chip, equipped with on-chip parameters and specialized modules including Padding, Dataflow Orchestrator, Dual-side Sparsity Detector, and Maxpooling.
Figure 5: Inter-layer Pipelined Architecture of FireFly-S. All cores are mapped onto a single FPGA chip, with each core acting as a pipeline stage responsible for the inference of one layer. These stages integrate modules for Padding, Dataflow Orchestration, Dual-side Sparsity Detection, and Maxpooling.
The Padding module dynamically adjusts input dimensions for convolutional layers, adding padding in ”same” mode to match output and input sizes, while bypassing padding for ”valid” mode or fully connected layers. The Dataflow Orchestrator is pivotal in ensuring consistent dataflow between layers to maintain pipeline efficiency. It also efficiently manages the im2col operation for SNNs, essential for reorganizing input data into a format suitable for convolution operations, all while maintaining a membrane potential storage-free approach. The Dual-side Sparsity Detector module optimizes computations by identifying matching nonzero elements in both activation spikes and synaptic weights, enabling computational savings by bypassing unnecessary operations. Then, it generates specific read commands for relevant data (weights, biases, etc.) and efficiently updates neuron dynamics in the IF/LIF neuron model. Finally, the optional Maxpooling module provides spatial downsampling of the output, the output stream is then pipelined and fed into the next layer for further processing.
@@ V-A Dataflow Orchestrator
To reconcile these conflicting dataflow requirements and manage the im2col operation, FireFly-S employs a dedicated Dataflow Orchestrator, as illustrated in Fig. 6. This FIFO interface logic utilizes on-chip RAM divided into two sectors. The first sector is essential for three key functions: 1) dataflow permutation across the temporal and channel dimensions, 2) execution of the implicit im2col operation, and 3) data reuse for continuous operation when parallelism cannot fully unroll the convolution. The size of this sector is determined by the input dimensions and kernel size, allowing it to store only the essential data needed to maintain the convolution window while the input feature map streams in. Additionally, both dataflow permutation and data reuse occur within the convolution window, with the reuse time calculated as $C_{o}/P_{C_{o}}$. The remaining on-chip RAM serves as a configurable holding area, acting as a buffer for incoming spike feature maps from upstream processing. This holding area prevents pipeline stalls that would otherwise occur when the reused sector is full. Additionally, it provides a transition buffer for switching convolution windows. For clarity, the example in Fig. 6 depicts a parallelism level of 1, with an input dataflow of $(F_{h_{i}},F_{w_{i}},C_{i},T)$ being transformed to the output dataflow of $(F_{h_{i}},F_{w_{i}},C_{o},T,K_{h},K_{w},C_{i})$.
@@
The orchestrator leverages a combination of status registers, implemented as simple counters, and some stride counters. The status registers, chained to form loop iterators, pinpoint the current position within the dataflow. These registers don’t directly determine memory addresses but provide signals that guide the stride counters in generating the correct addresses for fetching data from the RAM. The use of stride counters for base address and offset address, each with its own specific increment value, allows the orchestrator to navigate the non-sequential memory access patterns required by the implicit im2col, data transpositions and data reuse, eliminating the need for computationally expensive multiplication operations.
@@
The orchestrator manages data access using a push pointer and a pop pointer, reflecting the principles of a traditional FIFO buffer. While the logic for the push pointer aligns with standard FIFO operations, the pop pointer is determined by the combined values of the base and offset counters. The base counter indicates the starting position of the current convolution window within the RAM, while the offset counter navigates through the elements within that window.
@@
For valid convolution with stride 1, the base counter increment for the dataflow coordinator adapts based on the window’s position within the feature map. Under normal operation, the increment is simply $C_{i}\times T$, reflecting the number of data elements in the previous window’s channel and temporal dimensions. In addition to this regular increment, specific adjustments are needed for efficient transitions between rows and feature maps. After processing the last window of a row, the base counter must jump to the beginning of the next row, requiring an increment of $(K_{w}\times C_{i}\times T)$ to bypass the remaining $(K_{w}-1)$ columns of the previous row. Similarly, transitioning to a new feature map necessitates an increment of $((K_{h}-1)\times F_{w_{i}}+K_{w})\times C_{i}\times T$ to address the start of the new feature map’s data. This ability to directly access data in RAM, regardless of position, boosts the efficiency of the Dataflow Orchestrator. In contrast, traditional line buffers, which rely on sequential shifting, introduce overhead during row and feature map transitions, as they must shift data sequentially to align with the required starting position for processing.
@@
After processing the convolution window, the offset counter resets to its initial value, while the base counter remains unchanged during data reuse. Once the reuse is complete, the base counter increments to indicate the starting position of the next convolution window, as previously depicted. This cyclical process ensures that all output channels $C_{o}$ are effectively traversed while reusing the input data within the constraints of limited $P_{C_{o}}$ parallelism.
Besides its role in managing the dataflow, the design logic of Dataflow Orchestrator also leverages the regularity of data access patterns, which are dictated by consistent stride patterns, to pre-calculate base and offset counter values. This pre-calculation enables pipelined address generation for the pop pointer, efficiently determining the next data element for processing. The predictable nature of these strides also facilitates the implementation of the Dataflow Orchestrator’s full and empty signals as registers, reducing combinational logic delays and enhancing timing closure. Furthermore, as data dimensions increase, the bit width of the status and address registers scales logarithmically. This scaling ensures that the logic remains efficient and does not become a bottleneck during network expansion. Ultimately, these design choices support a working frequency of 333 MHz across our entire architecture, as detailed in Section VI.
[…]
@@ V-B Dual-side Sparsity Detector
The lower left corner of Fig. 7 depicts the initial transformation of the original sparse weight matrix into Bitmap format. This transformation, conducted offline before hardware deployment, evaluates each matrix element to determine its non-zero status, resulting in the creation of a bitmap mask alongside a separate collection of non-zero values. These bitmaps are then segmented into vectors, each sized according to the input channel parallelism, $P_{C_{i}}$. This segmentation, also performed pre-deployment based on workload estimations detailed in Section VI-B, ensures balanced processing across layers to mitigate pipeline stalls.
@@ VI-A Experiments Setup
To evaluate the performance of FireFly-S, we conduct a series of experiments using various SNN models, specifically SCNN5, SCNN7, and SCNN9 models with 5, 7, and 9 convolutional layers. Maintaining consistency with the previous FireFly iteration, we utilized LIF neurons in all experiments, with a time step of 4. Detailed configurations of these models are presented in Table II, where ’c3p1’ indicates a 3x3 convolution kernel with padding of 1, ’mp2’ represents max pooling with a 2x2 filter and stride of 2, and ’10fc’ represents a fully connected layer with 10 output neurons. These models are trained on the MNIST, DVS-Gesture, and CIFAR-10 datasets, respectively. For testing, all models are deployed on the Xilinx Zynq Ultrascale+ FPGA, utilizing the KV260 Vision AI Starter Kit, which is specifically designed for the deployment of edge-based applications.